A line has x-intercept a^2 and y-intercept ab (a and b are positive). Show that the equation of this line can be written as x/a + y/b = a
Answer by ewatrrr(24785)

 
Hi
x-intercept a^2 and y-intercept ab
(a^2,0)
(0, ab) m = -ab/a^2 = -b/a
y = (-b/a)x + ab
y/b = -x/a + a
x/a + y/b = a